OpenStudy (anonymous):

The math club and the science club had fundraisers to buy supplies for a hospice. The math club spent $135 buying six cases of juice and one case of bottled water. The science club spent $110 buying four cases of juice and two cases of bottled water. How much did a case of juice cost? How much did a case of bottled water cost?

OpenStudy (anonymous):

You set up the systems of equations like this: 135= 6j + 1w 110= 4j +2w Then you multiply the top equation by 2 (just multiply both sides by 2) 270=12j+2w Subtract the bottom equation from this 270-110= (12j+2w)- (4J+2w) 160=8j j=20 Then substitute back in to solve for the water. 110= 4(20) + 2w 30= 2w w=15 The price for a case of bottled water was 15$ and the price for a case of juice was 20$ Hope this helps!